OpenStudy (anonymous):

12r^4s^2+6r^3s^2-6r^2s^2 can you please show the solution ? :)

OpenStudy (saifoo.khan):

\[6 r^2 s^2 (r+1) (2 r-1)\]

OpenStudy (anonymous):

can you pls show the solution?

OpenStudy (saifoo.khan):

now, just solve this for zero..

OpenStudy (saifoo.khan):

6r^2s^2 = 0 r+1 = 0 2r-1 = 0

OpenStudy (anonymous):

tnx